Express the negation of each of these statements in terms




of quantifiers without using the negation symbol.




a) ∀x(x > 1)




b) ∀x(x ≤ 2)




c) ∃x(x ≥ 4)




d) ∃x(x < 0)




e) ∀x((x < −1) ∨ (x > 2))




f ) ∃x((x < 4) ∨ (x > 7))

1
Expert's answer
2022-05-30T23:36:43-0400

In general:

1) The negation of "\forall x (P(x))" is "\exist x(\lnot P(x))."

2) The negation of "\exist x (P(x))" is "\forall x (\lnot P(x))."


So we have:

a) The negation of "\u2200x(x > 1)" is "\exist x (\lnot (x>1))=\exist x(x \leq 1)."

b) The negation of "\u2200x(x \u2264 2)" is "\exist x (\lnot(x ≤ 2)) = \exist x (x>2)."

c) The negation of "\u2203x(x \u2265 4)" is "\forall x (\lnot(x ≥ 4)) = \forall x(x<4)."

d) The negation of "\u2203x(x < 0)" is "\forall x(\lnot(x < 0)) = \forall x (x ≥0)."

e) The negation of "\u2200x((x < \u22121) \u2228 (x > 2))" is "\exist x (\lnot ((x < −1) ∨ (x > 2))) = \exists x ((x ≥ -1) \land (x \leq 2))="

"\exist x (-1 \leq x \leq 2)."

f) The negation of "\u2203x((x < 4) \u2228 (x > 7))" is

"\forall x (\neg((x < 4) ∨ (x > 7))) = \forall x ((x ≥ 4) \land (x \leq 7))="

"=\forall x (4 \leq x \leq 7)."